If a simple pendulum oscillates with amplitude of 50mm and time period of 2s, then its maximum velocity is..
1 : 0.10 m/s
2 : 0.16 m/s
3 : 0.24 m/s
4 : 0.32 m/s
1
Expert's answer
2020-08-02T15:07:50-0400

Relationship of velocity and displacement 

"v=\omega\sqrt{A^2-x^2}"


"V=A\omega=A\frac{2\pi}{T}"


"50\times10^{-3}\times\frac{2\pi}{2}"


"V=50\times10^{-3}\times3.14"


"=0.16m\/s"

Therefore, option 2 : 0.16 m/s is correct